OpenStudy (anonymous):

@mathstudent55 :/ May you walk me through this one ..

OpenStudy (anonymous):

Okay... What you're going to want to do is to try and find the greatest common factor withing each row.

OpenStudy (anonymous):

In the case of the first row, what do 14, 21, 28 and 25 all have in common?

OpenStudy (anonymous):

35... my apologies

OpenStudy (anonymous):

just find the ratio of the corresponding sides of both figures given and see if the ratio is same for all corresponding pairs..

OpenStudy (anonymous):

7... and 2nd row is 8.. I don't understand how this relates to GCF...

OpenStudy (anonymous):

okay, so what are the factors that yield the numbers in each row? Are they the same in each row with the exception of the 7 and 8?

OpenStudy (anonymous):

like 16/14 =8/7 and 24/21 =8/7 if figures are similar their corresponding sides are proportional..

OpenStudy (anonymous):

you know what... listen to @matricked his way of doing this will probably work a bit more smoothly.

OpenStudy (mathstudent55):

Below are the 4 pairs of corresponding dimensions of the two pyramids. 16/14 24/21 32/28 40/35 Either reduce all the fractions above, or take a calculator and do all the divisions. What results do you get?

OpenStudy (anonymous):

1.14 1.14 ^ same again and again..

OpenStudy (mathstudent55):

That means all ratios are the same. When two solids have all corresponding linear dimensions in the same ratio, are they congruent or similar?

OpenStudy (anonymous):

They are similar, corresponding sides are the same; it is B!

OpenStudy (anonymous):

yup

OpenStudy (mathstudent55):

They are similar bec all pairs of corresp linear dimensions are in the same ratio You are correct..